Bologna sells for $1.50 a pound. Salami sells for $2.50 a pound. How much of each should be used to have ten pounds of a mixture of both selling for $1.75 a pound?
Answer by checkley75(3666)

2.50x+1.50(10-x)=10*1.75
2.50x+15-1.5x=17.50
x=17.50-15
x=2.50 pounds of salami
10-2.5=7.5 pounds of bologna
proof
2.50*2.5+1.50*7.5=17.50
6.25+11.25=17.50
17.50=17.50